Operations Report – April 17th

Crew 192 Operations Report 17Apr2018


Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 730 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133616.7

Toilet tank emptied: No. Emptied yesterday, April 16.

ATV’s Used: Red ATV, Blue ATV’s #1, #2, and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All ATVs working well, oil checked, gas filled today.

Deimos rover used: No
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31.2 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 91%
Ending charge: 91%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, trying to make the best use of time for EVAs in the midst of constant strong winds.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three two-piece suits used, radios working fine. Suit #4 charging port is faulty.

Summary of Hab operations: Cleaned and organized hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water was cleaned for final time today, remains in the RAM while strong winds persist.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Doctor Narayanamoorthi recommends biological and chemical testing of water in trailer water tank before clearance for crew use.

Faucet in upper deck of hab was fixed as best we could but it continues to slowly leak.

Copyright © The Mars Society. All rights reserved. | Main Site